OpenStudy (anonymous):

Jeff and Lucy have been asked to wash their mom’s minivan. It takes Jeff 2 hours to wash the van by himself, and it takes Lucy 1.5 hours to wash the van by herself. How long will it take Jeff and Lucy to wash the van if they work together? A.What equation is used to solve this problem? What does each variable represent? B.Identify the given information. C.Enter the given information into the equation. D.Solve the equation. E.State the solution.

OpenStudy (anonymous):

Let x be the number of hours for both to wash the car. What does1/x represent?

OpenStudy (anonymous):

1/x represents what portion of the car is washed by both them in one hour.

OpenStudy (anonymous):

Hence \[ \frac 1 x = \frac 1{1.5} + \frac 1{2} \]

OpenStudy (anonymous):

Solve for x

OpenStudy (anonymous):

Okay so what you basically need to do is 1/2+1/1.5=7/6, and then you can reciprocate it, so the answer would be 6/7

OpenStudy (anonymous):

You find x=6/7

OpenStudy (anonymous):

thank you all........